# Physics-Informed Dynamical Modeling of Extrusion-Based 3D Printing Processes

### v1

@@ 1 Introduction
Extrusion-based additive manufacturing (AM), particularly direct ink writing (DIW), has emerged as a transformative approach in the fabrication of cementitious structures for civil infrastructure, architectural applications, and rapid construction [1]. This process involves the layer-by-layer deposition of a viscoplastic cement-based ink, offering geometric freedom and reduced material waste. Despite significant advancements in hardware and material formulations, the fundamental understanding of the real-time monitoring and control of the underlying flow dynamics remain limited, particularly under varying geometric, material, and operational conditions.
@@
Traditional modeling approaches, rooted in continuum mechanics and finite element methods, offer detailed insights into localized flow and stress distributions but are computationally expensive and sensitive to complex rheological behavior. Moreover, these models often require precise knowledge of boundary conditions and material properties, which may be difficult to obtain or change during the printing process. These approaches convert partial differential equations (PDEs) to systems of algebraic equations that are solvable for computers. Examples of these approaches are the finite element method (FEM) and the finite volume method (FVM) [2]. There are multiple efforts in the existing literature that focus on modeling the flow of concrete using these simulation methods.
@@
A recent study by [3], focused on simulation-based flow modeling of five different non-Newtonian ink groups by numerically solving the conservation of mass and momentum equations. The results reveal the effect of the inlet pressure on the printability of the inks. However, in this study, the printability of the cement-based ink and the effect of the input parameters on the printed geometry have not been discussed. Some studies in the literature specifically studied the geometric accuracy and cross-sectional shape of the printed material. [4] analyzed the cross-sectional shape and dimensions of the printed bead by modeling the flow of the extruded material using FLOW-3D computational fluid dynamics (CFD) software. Different constitutive models were developed to study the shear-thinning and viscosity effects on the cross-section of the beads, and the results were compared with theoretical values used by the slicer software. To overcome the limitations of commercial software, including FLOW-3D, [5] combined OpenFOAM framework with the overInterDyMFoam solver to model 3D concrete printing to investigate the strand shape and cross-sectional geometry and dimensions in cases with various extrusion velocity, nozzle velocity and nozzle rotation. Similarly, [6] created a CFD model to simulate strand cross-sections at different nozzle gap height and velocity ratios. They experimentally validated the numerical model. Despite the high resolutions of the aforementioned studies, there is a contradiction between resolution and efficiency in modeling of additive manufacturing of cementitious materials using the DIW process. Moreover, the models developed by these simulation methods are computationally heavy, making them unsuitable for real-time applications.
@@
In response to these challenges, data-driven dynamical modeling has emerged as a promising framework for characterizing and predicting the behavior of extrusion-based 3D printing processes. By leveraging high-resolution spatiotemporal data from experiments or simulations, these models can capture complex, nonlinear system dynamics with reduced-order representations, enabling real-time monitoring, control, and optimization. The literature mostly focuses on the reduced-order control-oriented modeling of metal additive manufacturing techniques such as laser powder bed fusion ([7]) and directed energy deposition ([8]) to predict the geometrical characteristics of the printed part. However, there are only a few studies that focus on developing reduced-order models of the dynamics of the extrusion process.
@@
In this regard, in a study by [9], thermo-viscoelastic behaviors of the filaments in the extrusion-based fused deposition modeling process are captured by a reduced-order model. [10] developed models that can be used to understand spatial in-layer dynamics and layer-to-layer stability. [11] created a model of the filament behavior depending on the layer height. The model predicts the required applied force at the inlet depending on the layer height. Despite their efficiency and usefulness, these reduced-order models do not capture the evolution of the extruded strand as an explicit function of the governing process parameters. To the best of the authors’ knowledge, no existing study has formulated a reduced-order model specifically for the flow dynamics in DIW. This gap highlights the need for a computationally tractable yet physically meaningful model capable of predicting DIW strand formation based on controllable input variables.
@@ 2.2 Architecture of the dynamical model
In order to enable a tractable and simplified modeling approach, we visualize the input-output dynamics of the overall printing process as a cascade connection of three sub-systems, as shown in Fig. 2. The physical representations of these three sub-systems are shown in Fig. LABEL:fig:system1. First, we consider sub-system 1 where the inputs are the mass flow rate $\dot{m}$ of the material into the nozzle. The sub-system output is the flow velocity at the nozzle outlet. A simplified schematic of sub-system 1 is shown in Fig. 3.
@@
where $p_{1}$ is the pressure, $\rho$ is the density, $\nu$ is the kinematic viscosity, and $u_{1}$ and $v_{1}$ are the x and y velocity components, respectively. Here, the effective force ${f_{t}}_{1}$.
Furthermore, the velocity of the fluid is zero at the nozzle boundaries in the $x$-direction, due to the no-slip boundary condition:
@@
where the first term $\beta_{1}\,{p_{d}}_{1}$ is inspired by the term $\frac{1}{\rho}\frac{\partial p_{1}}{\partial y}$ in (14) with ${p_{d}}_{1}$ being the pressure gradient, and the second term $\beta_{2}\,\bar{v}_{1}$ is inspired by the term $3\nu\frac{\bar{v}_{1}}{m^{2}}$ in (14). The third term $\beta_{3}\,\dot{m}$ is inspired by the term $\frac{1}{2m}\int_{-m}^{m}{f_{t}}_{1}\,dx$ – where it is assumed that the major contribution of the force ${f_{t}}_{1}$ comes from the input mass flow rate $\dot{m}$. Here, the parameters $\beta_{1},\beta_{2}$ and $\beta_{3}$ are the tuning parameters of the model which need be fitted to match the ground truth data. This concludes the reduced order modeling of sub-system 1, which captures the dynamic relationship between the inlet mass flow rate $\dot{m}$ and pressure gradient ${p_{d}}_{1}$ with the output average velocity $\bar{v}_{1}$.
Figure 4: Schematic of sub-systems 2 and 3.
@@
The governing equations for sub-system 3 are similar to (1) and (2), with the following modifications: (i) The velocities are denoted as $u_{3}$ and $v_{3}$, instead of $u_{1}$ and $v_{1}$. (ii) The forcing term is dominated by the pressure gradient which now depends on the sub-system 2 outlet velocity $\bar{u}_{2}$.
@@
The flow at $y=0$ is attached to the moving surface and due to the no-slip condition, the flow velocity at $y=0$ is equal to the velocity of moving surface, $U_{s}$. At $y=h$ Free-surface boundary condition is applied. The resulting boundary conditions are:
@@
The governing equation (17) along with the boundary conditions (18), gives rise to a parabolic velocity profile ([12]). Accordingly, we choose a second-degree polynomial to capture the velocity profile:
@@ 2.3 Input-dependent model parameterization
Here, we assumed that sub-system 1 velocity dynamics is dependent on the inlet mass flow rate since this sub-system is directly affected by the inlet material injection, and the underlying nonlinearities are captured by (28) where the model coefficients are expressed as a linear function of $\dot{m}$. Similarly, the sub-system 3 velocity dynamics is dependent on the plate velocity since this sub-system is in direct contact of the plate, and the underlying nonlinearities are captured by (29) where the model coefficients are expressed as a linear function of $U_{s}$.
@@ 3 Data Generation and Reduced-Order Model Fitting
3.1 Data generation from high fidelity computational model
High-fidelity simulation models are commonly employed to evaluate the performance of the proposed model. CFD enables the numerical solution and analysis of the Navier–Stokes equations without requiring major simplifying assumptions. In this study, velocity and pressure fields within the three subsystems of the extrusion process are obtained from a 3D CFD simulation executed on an Intel Xeon® Gold 6230R CPU @ 2.10 GHz workstation with 128 GB of RAM. ANSYS Fluent (Workbench 2022R1) is used to simulate the flow of a cement-based material through the extrusion nozzle and its subsequent deposition onto a build plate. A pressure-based, transient solver with absolute velocity formulation is employed. Pressure–velocity coupling is achieved using the PISO algorithm, with both skewness and neighbor correction parameters set to 1. The PRESTO scheme is applied for pressure interpolation, second-order upwind for momentum, and compressive for volume-fraction interpolation. The least-squares cell-based method is used for gradient reconstruction, and a relaxation factor of 0.75 is applied to the flow variables. The CFD model has been validated against available literature results ([13]), showing good agreement in deposition profile. A detailed discussion of the validation and comparison with experimental results will be presented in a forthcoming publication.
@@ 3.1 Data generation from high fidelity computational model
3.2 Training and testing of the proposed dynamical model
In this subsection, we utilize the data generated from simulation cases in Table 1 to train the proposed dynamical model (given by (15), (16), and (27)) and test its performance. The ultimate objective of model training is to tune the model parameters (i.e., $\beta_{4}$ in (16) and $\gamma$’s in (28) and (29)) to minimize the loss between the model’s output and the ground-truth data generated by the CFD model. We have utilized the nonlinear least squares technique to fit the model parameters in MATLAB [14]. In the nonlinear least squares setting, the following objective function is minimized
$\min_{x}f(x)=\sum_{i=1}^{n}f_{i}(x)^{2},$ (30)
in terms of auxiliary functions ($f_{i}(x)$) and the algorithm minimizes the sum of squares of these functions. Suppose that the observed-data from the CFD simulation is in the form of $\{(t_{i},y_{i})\}_{i=1}^{n}$ and the model is in the form of $\hat{M}(\gamma)$, then the algorithm solves the following equation to estimate parameter $\gamma$:
Hence, the model $\hat{M}$ is essentially a time-discretized version of the model (15), (16), and (27) (using Euler’s first order discretization) and the parameters are $\beta_{4}$ and $\gamma_{j}$ where $j\in\{1,2,3,5,6,7\}$.
@@ 4.1 High-fidelity data analysis
The inlet velocity is prescribed at three discrete values, 50 mm/s, 60 mm/s, and 70 mm/s, corresponding to inlet mass flow rates of 0.02475 kg/s, 0.02970 kg/s, and 0.03465 kg/s, respectively. Independently, the surface velocity is varied across the same three levels (50 mm/s, 60 mm/s, 70 mm/s) to emulate different nozzle–substrate relative motion speeds during deposition. These values are selected based on the prescribed values in the literature for the DIW of the cement-based materials [13]. A full factorial combination of these two parameters results in the nine simulation cases summarized in Table 1, ensuring that both primary process inputs are uniformly sampled across the design space.
@@
Taken together, these observations provide a consistent and mechanistic understanding of flow behavior across the DIW process. The nozzle interior (sub-system 1) operates in a pressure-dominated regime where inlet mass flow rate is the primary driver. The downstream region on the build plate (sub-system 3) operates in a kinematically driven regime where substrate motion dominates. The intermediate region (sub-system 2) embodies a coupled regime in which both inputs shape the velocity field, reflecting the interplay between extrusion-induced flow acceleration, strand formation, and substrate-imposed shear. This physics-based interpretation forms the foundation for subsequent model training and validation across varying process conditions.
@@ 4.2 Model evaluation for different mass flow conditions
In this subsection, first, we assess the ability of the dynamical model to perform interpolation with respect to inlet mass flow rate. To this end, the model is trained using datasets corresponding to the lowest and highest mass flow rates in the simulation set, specifically 0.02475 kg/s and 0.03465 kg/s. After training, the model is evaluated on the intermediate mass flow rate of 0.02970 kg/s, a condition not included in the training set. This testing configuration enables investigation of the model’s capability to generalize and accurately predict flow dynamics for mass flow rates that lie between the trained extremes.
@@
Figure 7: Performance of the dynamical model trained with simulation cases 1, 2, 3, 7, 8, and 9. The model is evaluated on intermediate flow conditions (cases 4, 5, and 6) not used during training. “Data” indicated CFD simulation data and “Model” indicates proposed dynamical model. The $y$-axis for the first column corresponds to $\bar{v}_{1}$ (mm/s), the second column to $\bar{u}_{2}$ (mm/s), and the third column to $\bar{u}_{3}$ (mm/s), while the $x$-axis for all figures represents time in seconds (s).
@@
For sub-system 3, the model once again produces error distributions that are tightly clustered around zero and generally bounded within 0.005-0.005 m/s to $0.005$ m/s, with some data points extending beyond $0.005$ m/s. This consistent error confinement across both training and testing scenarios demonstrates the robustness of the model in capturing the corresponding velocity profiles, particularly under standard operating conditions where flow features may be less complex. Overall, the results highlight the model’s strong predictive capability across the three sub-systems, with especially high accuracy observed for sub-systems 1 and 3 and acceptable, well-bounded performance for sub-system 2.
@@
The RMSE distribution in Figure 10(a) illustrates the predictive performance of the model when trained on higher-end extrapolative mass flow rates. The results show improved prediction accuracy for all sub-systems compared to the interpolative training case, with sub-system 1 exhibiting particularly strong performance. This indicates that the model is capable of reliably extrapolating to slightly lower mass flow rates when trained on high-end data. In contrast, for sub-systems 2 and 3, the prediction errors are larger than both the training errors and those observed in the interpolative case. This increased error arises because these sub-systems are influenced by additional factors such as surface velocity and coupled flow effects that are not fully represented in the higher-end training data, limiting the model’s ability to extrapolate accurately in these regions.

### v2

@@ 1 Introduction
Extrusion-based additive manufacturing (AM), particularly direct ink writing (DIW), has emerged as a transformative approach in the fabrication of cementitious structures for civil infrastructure, architectural applications, and rapid construction [mechtcherine2020extrusion]. This process involves the layer-by-layer deposition of a viscoplastic cement-based ink, offering geometric freedom and reduced material waste. Despite significant advancements in hardware and material formulations, the fundamental understanding of the real-time monitoring and control of the underlying flow dynamics remain limited, particularly under varying geometric, material, and operational conditions.
@@
Traditional modeling approaches, rooted in continuum mechanics and finite element methods, offer detailed insights into localized flow and stress distributions but are computationally expensive and sensitive to complex rheological behavior. Moreover, these models often require precise knowledge of boundary conditions and material properties, which may be difficult to obtain or change during the printing process. These approaches convert partial differential equations (PDEs) to systems of algebraic equations that are solvable for computers. Examples of these approaches are the finite element method (FEM) and the finite volume method (FVM) [roussel2020numerical]. There are multiple efforts in the existing literature that focus on modeling the flow of concrete using these simulation methods.
@@
A recent study focused on simulation-based flow modeling of five different non-Newtonian ink groups by numerically solving the conservation of mass and momentum equations [sourov2025general]. The results reveal the effect of the inlet pressure on the printability of the inks. However, in this study, the printability of the cement-based ink and the effect of the input parameters on the printed geometry have not been discussed. Some studies in the literature specifically studied the geometric accuracy and cross-sectional shape of the printed material. For example, the cross-sectional shape and dimensions of the printed bead were analyzed by modeling the flow of the extruded material using FLOW-3D computational fluid dynamics (CFD) software [rapha]. Different constitutive models were developed to study the shear-thinning and viscosity effects on the cross-section of the beads, and the results were compared with theoretical values used by the slicer software. To overcome the limitations of commercial software, including FLOW-3D, OpenFOAM framework was combined with the overInterDyMFoam solver to model 3D concrete printing to investigate the strand shape and cross-sectional geometry and dimensions in cases with various extrusion velocity, nozzle velocity and nozzle rotation [wagner2025novel]. Similarly, an experimentally-validated CFD model was created to simulate strand cross-sections at different nozzle gap height and velocity ratios [serdeczny2018experimental]. Despite the high resolutions of the aforementioned studies, there is a contradiction between resolution and efficiency in modeling of additive manufacturing of cementitious materials using the DIW process. Moreover, the models developed by these simulation methods are computationally heavy, making them unsuitable for real-time applications.
@@
In response to these challenges, data-driven dynamical modeling has emerged as a promising framework for characterizing and predicting the behavior of extrusion-based 3D printing processes. By leveraging high-resolution spatiotemporal data from experiments or simulations, these models can capture complex, nonlinear system dynamics with reduced-order representations, enabling real-time monitoring, control, and optimization. The literature mostly focuses on the reduced-order control-oriented modeling of metal additive manufacturing techniques such as laser powder bed fusion [wang2019control] and directed energy deposition [wang2016reduced] to predict the geometrical characteristics of the printed part. However, there are only a few studies that focus on developing reduced-order models of the dynamics of the extrusion process.
@@
In this regard, in a recent study, thermo-viscoelastic behaviors of the filaments in the extrusion-based fused deposition modeling process were captured by a reduced-order model [meng2024reduced]. Advanced models were also developed to understand spatial in-layer dynamics and layer-to-layer stability [balta2019control]. In another study, the authors created a model of the filament behavior depending on the layer height [percoco2020analytical]. The model was able to predict the required applied force at the inlet depending on the layer height. Despite their efficiency and usefulness, these reduced-order models do not capture the evolution of the extruded strand as an explicit function of the governing process parameters. To the best of the authors’ knowledge, no existing study has formulated a reduced-order model specifically for the flow dynamics in DIW. This gap highlights the need for a computationally tractable yet physically meaningful model capable of predicting DIW strand formation based on controllable input variables.
@@ 2.2 Architecture of the dynamical model
In order to enable a tractable and simplified modeling approach, we visualize the input-output dynamics of the overall printing process as a cascade connection of three sub-systems, as shown in Fig. 2. The physical representations of these three sub-systems are shown in Fig. 1. First, we consider sub-system 1 where the input is the mass flow rate $\dot{m}$ of the material into the nozzle. The sub-system output is the flow velocity at the nozzle outlet. A simplified schematic of sub-system 1 is shown in Fig. 3.
@@
where $p_{1}$ is the pressure, $\rho$ is the density, $\nu$ is the kinematic viscosity, and $u_{1}$ and $v_{1}$ are the x and y velocity components, respectively. Here, the effective force is ${f_{t}}_{1}$.
Furthermore, the velocity of the fluid is zero at the nozzle side walls, due to the no-slip boundary condition:
@@
where the first term $\beta_{1}\,{p_{d}}_{1}$ is inspired by the term $\frac{1}{\rho}\frac{\partial p_{1}}{\partial y}$ in (14) with ${p_{d}}_{1}$ being the pressure gradient, and the second term $\beta_{2}\,\bar{v}_{1}$ is inspired by the term $3\nu\frac{\bar{v}_{1}}{m^{2}}$ in (14). The third term $\beta_{3}\,\dot{m}$ is inspired by the term $\frac{1}{2m}\int_{-m}^{m}{f_{t}}_{1}\,dx$ – where it is assumed that the major contribution of the force ${f_{t}}_{1}$ comes from the input mass flow rate $\dot{m}$. Here, the parameters $\beta_{1},\beta_{2}$ and $\beta_{3}$ are the tuning parameters of the model which need to be fitted to match the ground truth data. This concludes the reduced order modeling of sub-system 1, which captures the dynamic relationship between the inlet mass flow rate $\dot{m}$ and pressure gradient ${p_{d}}_{1}$ with the output average velocity $\bar{v}_{1}$.
@@
The governing equations for sub-system 3 are similar to (1) and (2), with the following modifications: (i) The velocities are denoted as $u_{3}$ and $v_{3}$, instead of $u_{1}$ and $v_{1}$. (ii) The forcing term is dominated by the pressure gradient, which now depends on the sub-system 2 outlet velocity $\bar{u}_{2}$.
@@
The flow at $y=0$ is attached to the moving surface and due to the no-slip condition, the flow velocity at $y=0$ is equal to the velocity of moving surface, $U_{s}$. At $y=h$, the free-surface boundary condition is applied. The resulting boundary conditions are:
@@
The governing equation (17) along with the boundary conditions (18), gives rise to a parabolic velocity profile ([samanta2024combined]). Accordingly, we choose a second-degree polynomial to capture the velocity profile:
@@ 2.3 Input-dependent model parameterization
Here, we assumed that the velocity dynamics of sub-system 1 depend on the inlet mass flow rate, since this sub-system is directly affected by material injection at the inlet, and the underlying nonlinearities are captured by (28), where the model coefficients are expressed as a linear function of $\dot{m}$. Similarly, the velocity dynamics of sub-system 3 depend on the plate velocity since this sub-system is in direct contact with the plate, and the underlying nonlinearities are captured by (29), where the model coefficients are expressed as a linear function of $U_{s}$.
@@ 3 Data Generation and Reduced-Order Model Fitting
3.1 Data generation from high-fidelity computational model
High-fidelity simulation models are commonly employed to evaluate the performance of the proposed model. CFD enables the numerical solution and analysis of the Navier–Stokes equations without requiring major simplifying assumptions. In this study, velocity and pressure fields within the three subsystems of the extrusion process are obtained from a 3D CFD simulation executed on an Intel Xeon® Gold 6230R CPU @ 2.10 GHz workstation with 128 GB of RAM. ANSYS Fluent (Workbench 2022R1) is used to simulate the flow of a cement-based material through the extrusion nozzle and its subsequent deposition onto a build plate. A pressure-based, transient solver with absolute velocity formulation is employed. Pressure–velocity coupling is achieved using the PISO algorithm, with both skewness and neighbor correction parameters set to 1. The PRESTO scheme is applied for pressure interpolation, second-order upwind for momentum, and compressive for volume-fraction interpolation. The least-squares cell-based method is used for gradient reconstruction, and a relaxation factor of 0.75 is applied to the flow variables. The CFD model has been validated against available literature results [liu2020modelling], showing good agreement in deposition profile. A detailed discussion of the validation and comparison with experimental results will be presented in a forthcoming publication.
@@ 3.1 Data generation from high-fidelity computational model
3.2 Training and testing of the proposed dynamical model
In this subsection, we utilize the data generated from simulation cases in Table 1 to train the proposed dynamical model (given by (15), (16), and (27)) and test its performance. The ultimate objective of model training is to tune the model parameters (i.e., $\beta_{4}$ in (16) and $\gamma$’s in (28) and (29)) to minimize the loss between the model’s output and the ground-truth data generated by the CFD model. We have utilized a nonlinear least-squares technique to fit the model parameters in MATLAB [dennis1977nonlinear]. In the nonlinear least-squares setting, the following objective function is minimized:
$\min_{x}f(x)=\sum_{i=1}^{n}f_{i}(x)^{2},$ (30)
where $f_{i}(x)$ are auxiliary functions, and the algorithm minimizes the sum of their squares. Suppose that the observed-data from the CFD simulation is in the form of $\{(t_{i},y_{i})\}_{i=1}^{n}$ and the model is in the form of $\hat{M}(\gamma)$, then the algorithm solves the following equation to estimate parameter $\gamma$:
Hence, the model $\hat{M}$ is essentially a time-discretized version of the model (15), (16), and (27) (using Euler’s first order discretization) and the parameters are $\beta_{4}$ and $\gamma_{j}$, where $j\in\{1,2,3,5,6,7\}$.
@@ 4.1 High-fidelity data analysis
The inlet velocity is prescribed at three discrete values, 50 mm/s, 60 mm/s, and 70 mm/s, corresponding to inlet mass flow rates of 0.02475 kg/s, 0.02970 kg/s, and 0.03465 kg/s, respectively. Independently, the surface velocity is varied across the same three levels (50 mm/s, 60 mm/s, 70 mm/s) to emulate different nozzle–substrate relative motion speeds during deposition. These values are selected based on the prescribed values in the literature for the DIW of the cement-based materials [liu2020modelling]. A full factorial combination of these two parameters results in the nine simulation cases summarized in Table 1, ensuring that both primary process inputs are uniformly sampled across the design space.
@@
Taken together, these observations provide a consistent and mechanistic understanding of flow behavior across the DIW process. The nozzle interior (sub-system 1) operates in a pressure-dominated regime, where the inlet mass flow rate is the primary driver. The downstream region on the build plate (sub-system 3) operates in a kinematically driven regime, where substrate motion dominates. The intermediate region (sub-system 2) embodies a coupled regime in which both inputs shape the velocity field, reflecting the interplay between extrusion-induced flow acceleration, strand formation, and substrate-imposed shear. This physics-based interpretation forms the foundation for subsequent model training and validation across varying process conditions.
@@ 4.2 Model evaluation for different mass flow conditions
In this subsection, we first assess the ability of the dynamical model to perform interpolation with respect to the inlet mass flow rate. To this end, the model is trained using datasets corresponding to the lowest and highest mass flow rates in the simulation set, specifically 0.02475 kg/s and 0.03465 kg/s. After training, the model is evaluated on the intermediate mass flow rate of 0.02970 kg/s, a condition not included in the training set. This testing configuration enables investigation of the model’s capability to generalize and accurately predict flow dynamics for mass flow rates that lie between the trained extremes.
@@
Figure 7: Performance of the dynamical model trained with simulation cases 1, 2, 3, 7, 8, and 9. The model is evaluated on intermediate flow conditions (cases 4, 5, and 6) not used during training. “Data” indicates CFD simulation data and “Model” indicates proposed dynamical model. The $y$-axis for the first column corresponds to $\bar{v}_{1}$ (mm/s), the second column to $\bar{u}_{2}$ (mm/s), and the third column to $\bar{u}_{3}$ (mm/s), while the $x$-axis for all figures represents time in seconds (s).
@@
For sub-system 3, the model once again produces error distributions that are tightly clustered around zero and generally bounded within 0.005-0.005 m/s to $0.005$ m/s, with some data points extending beyond $0.005$ m/s. This consistent error confinement across both training and testing scenarios demonstrates the robustness of the model in capturing the corresponding velocity profiles, particularly under standard operating conditions, where flow features may be less complex. Overall, the results highlight the model’s strong predictive capability across the three sub-systems, with especially high accuracy observed for sub-systems 1 and 3 and acceptable, well-bounded performance for sub-system 2.
@@
The RMSE distribution in Figure 10(a) illustrates the predictive performance of the model when trained on higher-end extrapolative mass flow rates. The results show improved prediction accuracy for all sub-systems compared to the interpolative training case, with sub-system 1 exhibiting particularly strong performance. This indicates that the model is capable of reliably extrapolating to slightly lower mass flow rates when trained on high-end data. In contrast, for sub-systems 2 and 3, the prediction errors are larger than both the training errors and those observed in the interpolative case. This increased error arises because these sub-systems are influenced by additional factors, such as surface velocity and coupled flow effects, that are not fully represented in the higher-end training data, limiting the model’s ability to extrapolate accurately in these regions.